OpenStudy (anonymous):

how to find this limit: lim (xˆ2-1)/(x-1) when x goes to 1?

OpenStudy (anonymous):

2 is the answer

OpenStudy (anonymous):

2 because:\[\frac{x^2-1}{x-1} = \frac{(x-1)(x+1)}{x-1} = x+1\]

OpenStudy (anonymous):

(x^2-1)/(x-1) (x-1)(x+1)/(x-1) (x+1) after applying limit x goes to 1 we have answer is 2